OpenStudy (anonymous):

write y=0.81x+25 in standard form

OpenStudy (mertsj):

.81x-y=-25

OpenStudy (anonymous):

-0.81x+y=25

Directrix (directrix):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. ------------------ .81x + (-1) y = -25